OpenStudy (anonymous):

2m - n = 15 2m + n = 3

hartnn (hartnn):

if you want to do it using substitution only, then find 2m from first equation and substitute it in 2nd equation.

OpenStudy (aravindg):

add the 2 equations .what do u get?

OpenStudy (anonymous):

2m-n= 15 ------(1) 2m+n=3 ------(2) Remove one variable first as the constant before n is 1 and -1, we can cancel the two n by adding (1) and (2) (1)+(2) : 4m=18 so m=9/2 Until this, there are two methods to continue our work. Method 1: sub n to any one of the equations. For example, sub. m=9/2 into (1), 2(9/2) -n=15 n=-6 Method 2: (1)-(2) to remove m. -2n=12 n=-6

hartnn (hartnn):

i think substitution needs to be used ...isn't it ?

OpenStudy (anonymous):

Either way i still keep getting 4m = 18...

hartnn (hartnn):

thats correct, so m=... ?

hartnn (hartnn):

divide by 4 on both sides.

OpenStudy (anonymous):

I don't know how to solve that one. /.\

OpenStudy (aravindg):

why not??

OpenStudy (aravindg):